OpenStudy (anonymous):

a particular angle has a measure of 42 degrees 30'36". convert this angle measure to decimal notation. ***fan and medal***

OpenStudy (anonymous):

1 minute = 1/60 degree 1 second = 1/60 second = 1/3600 degree Then we have 42 degrees + 30/60 degree + 36/3600 degrees. 30/60 = 0.5 36/3600 = 0.01 42 + 0.5 + 0.01 = 42.51